OpenStudy (anonymous):

Help please? 5/x-5 + 3/4 = 2/x-5

OpenStudy (anonymous):

\[is \it \frac{ 5 }{ x-5 }\]

OpenStudy (anonymous):

is it*

OpenStudy (anonymous):

or 5/x -5

OpenStudy (anonymous):

it is x=1 just don't know how to get it

OpenStudy (anonymous):

hmmmm....

OpenStudy (dtan5457):

I did a similar problem like this a few days ago, multiply the x-5 by x+5 to eliminate the fractions for a start.

OpenStudy (anonymous):

sorry still lost

OpenStudy (dtan5457):

@Directrix @freckles

OpenStudy (dtan5457):

I have to go, sorry. Hopefully you get more help, been procrastinating on my own work long enough...

OpenStudy (anonymous):

thanks!

OpenStudy (freckles):

\[\frac{5}{x-5}+\frac{3}{4}=\frac{2}{x-5} ?\]

OpenStudy (freckles):

look at the bottom of the fractions you (x-5) and (4) that are bottom factors multiply both sides by 4(x-5) to eliminate the fractions

OpenStudy (freckles):

that is \[4(x-5)\cdot [\frac{5}{x-5}+\frac{3}{4}]=\frac{2}{x-5} \cdot 4(x-5) \\ \text{ distribute on the left } \\ 4(x-5) \frac{5}{x-5}+4(x-5)\frac{3}{4}=\frac{2}{x-5}4(x-5)\] do some canceling and tell me your resulting equation

OpenStudy (anonymous):

20+3(x-5)=8?

OpenStudy (anonymous):

I get it now thank you!

OpenStudy (freckles):

and yes you got exactly the right equation

OpenStudy (freckles):

gj

OpenStudy (freckles):

you can solve that?

OpenStudy (anonymous):

yes thank u !